OpenStudy (baby456):

solve this systems of equations help asap medal + fan -2x+y=8 y=-3x+2

OpenStudy (solomonzelman):

-2x+y=8 --> y=2x+8 y=-3x+2 2x+8=-3x+2

OpenStudy (solomonzelman):

solve for x, and then find y.

OpenStudy (baby456):

opps its 3x-2 actually

OpenStudy (baby456):

i typed it wrong

OpenStudy (baby456):

x=2

OpenStudy (baby456):

how do i solve for y

OpenStudy (baby456):

okay sorry i was solving this in my head by i got cordinates 2,4 because it was supposed to be -2x+8=-3x-2 thank you

OpenStudy (baby456):

i got 2 from x values and got 4 by putting 2 into -2x+y=8 thankyou @SolomonZelman